A project has cash flows of -$152,000, $70,800, $62,500 and $95,000 for years 0 to 3, respectively. The required rate of return is 10 percent. What is the profitability index?
Answer #1

NPV = PV of Cash Flows - Investment =70800/(1+10%)+62500/(1+10%)^2+95000/(1+10%)^3-152000 =35391.44

Profitability Index =1+NPV/Investment =1+35391.44/152000 =1.23